A product has the following demand and supply equations:
Qd=100-3P and Qs=20+2P

a) What is the equilibrium price?
b) What is the equilibrium quantity?
1
Expert's answer
2020-05-29T09:34:40-0400

Equilibrium means that supply is equial to demand:

"Q_d=100-3P=20+2P=Q_s"

From this equation:

a)Equillibrium price is "P=16"

b)Equillibrium quantity is "Q_d=Q_s=20+2P=52"